SCI Question 1

SOLVE FOR THE NET INCOME

 

Gross Sales = 5,000,000 pesos

Sales Discounts = 230,000 pesos

Cost of Goods Sold= 3,120,000 pesos

Distribution Expenses= 500,000 pesos

Administrative Expenses= 120,000 pesos

Income Tax Expense= 25% of Operating Profit


1
Expert's answer
2021-09-27T15:45:22-0400

Net Income = Gross Profit - all other expenses

Gross Profit = Revenue - Cost of Goods Sold

Gross Sale - Sales Discount = Net Sales

= 5,000,000 - 230,000 =4,770,000


Gross Profit = 4,770,000 - 3,120,000 = 1,650,000 pesos


Operating Profit = Gross Profit - Operating Expenses

=Gross Profit - Distribution Expenses - Administration Expenses

= 1,650,000 - 500,000 - 120,000 = 1,030,000 pesos

Income Tax Expense = 25% × 1,030,000 = 257500 pesos


Net Income = Gross Profit - all other expenses

= Gross Profit - Distribution Expenses - Administration Expenses - Income Tax Expense

= 1650000 - 500000 -120000 - 257500 = 772500 pesos